OpenStudy (anonymous):

PLEASE HELP I WILL MEDAL AND FAN!!! Find the unit rate, and then use it to solve the problem. A gardening company plants 10 trees in 2 hours. How long will it take to plant 150 trees? A. 20 h B. 30 h C. 50 h D. 75 h

OpenStudy (welshfella):

10 trees in 2 hours so he plants 1 tree in 2/10 = 1/5 hours for 150 trees its 150 * 1/5 hours

OpenStudy (anonymous):

So the answer would be 30?

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Great Thank you for your help. : )

OpenStudy (welshfella):

yw

OpenStudy (texaschic101):

this could also be solved with a proportion...although the problem does state " solve for unit rate "...but if you didn't have to solve for unit rate. it could be solved with this : 10/2 = 150/x

OpenStudy (anonymous):

Cool

OpenStudy (anonymous):

I did not know that (mostly because I am terrible at math) but yeah cool. : )